Month Advertising Expenditures (x) Units Sold (y)
1 $2,215 543
2 $2,975 664
3 $2,150 538
4 $2,060 575

9. Using a 90% confidence level, does a change in advertising spending have a statistically significant impact on sales?

10. What is the correct interpretation of the coefficient of determination?

11. What are expected sales if advertising spending is $3,000?

12. What is the size of the error term for month #1?

13. Because the OLS method was used, what can be said about the fit of this line to the scatterplot of data?

14. The degrees of freedom for this model is:
Answer #1

The value of R is 0.9845.

This is a strong positive correlation, which means that high X variable scores go with high Y variable scores (and vice versa).

Click here to calculate a p-value.

9)9. Using a 90% confidence level, does a change in advertising spending have a statistically significant impact on sales?

The P-Value is .0308. The result is significant at p < .10.

10) The value of R2, the coefficient of determination, is 0.9692.

strong positive relation between Expenditures and sold

11)11. What are expected sales if advertising spending is $3,000?

3000=1474.1727+6.5934X

X=(3000-1474.1727 ) / 6.5934

X=231.1893

12)error term

Sum of squares (SSX) = 10214

MSE=10214/4

MSE=2553.5

13)

For your data, the regression equation for Y is:

ŷ = 6.5934X - 1474.1727

14. The degrees of freedom for this model is: 4